Need help with the following problem: only one point (x,y)

The next two parts deal with the potential away fr
Answer #1

(1, 7)

We require the point on the line y=7, such that the ratio of its distance from the positive charge to its distance from the negative charge is 1/3.

\therefore\ \frac{r_p}{r_n}\ =\ \frac{1}{3}

Let the point be at a distance x in the x-direction from the positive charge.

\therefore\ \frac{r_p}{r_n}\ =\ \frac{\sqrt{3^2\ +\ x^2}}{\sqrt{3^2\ +\ (8-x)^2}}\ =\ \frac{1}{3}

Solving for x, we get,

x = -1.

This means that the point has the x-coordinate at a distance of 1 unit towards the left of the positive charge.

Therefore, the co-ordinates of the equipotential point along y=7 are:

(1, 7)